OpenStudy (anonymous):

express the complex number in polar form there are two different problems 5.2i 6+4i

OpenStudy (anonymous):

(5.2) < 90 7.2111 < 33.69

OpenStudy (anonymous):

how tho i need to learn this :(

OpenStudy (anonymous):

\[i = cos(\frac{\pi}{2})+ i sin(\frac{\pi}{2})\] so first answer is \[5.2(cos(\frac{\pi}{2})+ i sin(\frac{\pi}{2}))\]

OpenStudy (anonymous):

great website to learn: http://www.intmath.com/complex-numbers/4-polar-form.php

OpenStudy (anonymous):

in general \[a+bi=r(cos(\theta)+isin(\theta))\] \[r=\sqrt{(a^2+b^2)}\] \[tan(\theta)=\frac{b}{a}\]

OpenStudy (anonymous):

perhaps i read the problem wrong. is the first one 5.2i or 5+2i?